# Roundtable Policy: Improving Scientific Reasoning and Narratives through Confidence-Weighted Consensus of LLMs

### v1

@@ C.1 Dataset Annotation
We introduce MatSci–Bench, a heterogeneous benchmark designed to evaluate the scientific reasoning capabilities of large language models (LLMs) in materials science and beyond. The benchmark spans six carefully curated sub-domains: (i) Macro Mechanics, (ii) Micro Mechanics, (iii) Optical & Microscopic, (iv) Semiconductor Physics, (v) Thermodynamical Systems, and (vi) Electrical & Electronic Properties. As visualized in Figure 8, these categories are balanced in both question count and token length, enabling systematic cross-domain evaluation of diverse reasoning skills.
@@
This hierarchical structure compels models to integrate reasoning across physical length scales—from atomic-level diffusion and microstructure evolution to bulk elasticity and device performance—while combining symbolic, conceptual, and numerical reasoning. Each sub-domain is paired with task-specific rubrics assessing both (a) factual accuracy and (b) explanatory depth. To further capture reasoning quality, all model outputs must include a chain-of-thought rationale, which is evaluated through two complementary methods: (i) the automatic Causal Reasoning Index (CRI), and (ii) a human-curated explanation rubric (details in Appendix C.4). These fine-grained metrics allow for robust measurement of model performance and principled ablation studies. Table 4 provides representative examples across four broader scientific areas—Biology, Geoscience, Mathematics, and Materials Science—highlighting domain-specific instruction design and diverse task objectives. Unlike prior benchmarks focused narrowly on short factual queries, our dataset emphasizes heterogeneous scientific tasks requiring numerical derivation, mechanistic explanation, and symbolic fluency.
Materials-science subset.
@@
The materials-science questions are drawn from the 2006–2024 UCLA Materials Science and Engineering Ph.D. qualifying examinations. This provenance ensures originality, rigor, and relevance to authentic scientific reasoning. Each problem is annotated with: (i) a numeric ground-truth answer rounded to three significant digits, (ii) an exact LaTeX expression preserving symbolic fidelity, and (iii) a step-by-step expert-verified solution. All annotations are fully compliant with LaTeX and Markdown standards, ensuring reproducibility and integration with downstream evaluation pipelines.
@@
Subdomain Task Instruction Subtask Math Task 007: Math Janet’s ducks lay 16 eggs per day. She consumes 3 eggs for breakfast, uses 4 eggs for baking muffins, and sells the remaining eggs at $2 each. How much does she earn at the farmers’ market? Display strong arithmetic and logical sequencing skills to compute the final earnings. Geoscience Task 008: Geo. Archaeologists record the top layers first, followed by older, deeper layers to study ancient settlements according to the principle of superposition. Demonstrate the ability to interpret stratigraphic evidence and correctly apply the principle of superposition. Biology Task 009: Bio. Does tranexamic acid reduce desmopressin-induced hyperfibrinolysis? Link pharmacological interventions with clinical outcomes based on empirical data. Materials Science Macro Mechanics Task 001: Mech. 1 What is elasticity? What is strain energy? How do we make use of high elastic modulus materials? Explain fundamental mechanical properties and their practical applications in structural design. Micro Mechanics Task 002: Mech. 2 What are interstitial and substitutional diffusion? How do they differ from self-diffusion, and what drives these processes? Differentiate between diffusion mechanisms and explain the role of chemical potential gradients. Electrical & Electronic Task 003: EE What is the difference between homogeneous and heterogeneous nucleation? Demonstrate understanding of nucleation processes and how surfaces or impurities influence phase transitions. Optical & Microscopic Task 004: Opt. Using characterization techniques, how would you distinguish between a mixture of 90% NaCl + 10% LiCl and a mixture of 90% NaCl + 5% LiCl + 5% KCl? Interpret diffraction patterns or elemental analysis data to differentiate material compositions. Thermodynamical Task 005: Thermo. Explain the laws of thermodynamics both mathematically and conceptually. Articulate both the formal mathematical expressions and the underlying concepts of the thermodynamic laws. Semiconductor Task 006: Semi. What is Schrödinger’s equation? Use it to solve for a 1D wavefunction in a simple potential. Demonstrate a solid understanding of quantum mechanics by providing the correct form of Schrödinger’s equation and a step-by-step solution for a 1D potential problem.
Success Criteria: The arbitrator agent confirms that the confidence score of the multi-agents converges.
@@ C.2 Dataset Comparison
| Geoscience | 1395 | 109,159 |
| Biology | 1000 | 92,866 |
| Math | 1319 | 246,238 |
| Material Science | 539 | 46,150 |
Table 5 benchmarks MatSci–Bench against a range of prominent scientific QA datasets. Whereas prior benchmarks often emphasize scale—providing thousands of short, loosely categorized questions—MatSci–Bench is designed for conceptual depth, compositional reasoning, and graduate-level rigor. For instance, the materials science subset averages tokens per question, more than double that of typical mathematics datasets, reflecting the structural and semantic complexity encountered in authentic scientific workflows.
@@ C.3 QK-Pair Overview
MultiTask. Each round presents a batch of heterogeneous tasks spanning multiple scientific domains as shown in Figure 9. The AI Committee jointly evaluates player responses, testing the arbitration system’s ability to generalize across diverse input types, modalities, and difficulty levels.
SingleTask. Each round focuses on a scientific proposal generation as shown in Figure 10, allowing the Committee to learn nuanced arbitration strategies within a fixed domain and structured output format.
Together, these protocols provide a comprehensive evaluation of arbitration dynamics under both breadth-oriented (MultiTask) and depth-oriented (SingleTask) conditions. Importantly, all evaluations are conducted without modifying the internal parameters of the player models, thereby preserving their pretrained competencies while enabling compositional reasoning through learned arbitration.
C.4 MatSci–Bench
Table 6 illustrates the mapping between several UCLA courses (from the prelim exams spanning 2006–2024) and our six defined subdomains in Material Science. The table indicates the number of questions per course and marks (✓) the corresponding subdomains covered by each course. And Table 7 Mapping from numeric course codes to official course names in the UCLA MSE Graduate Program.
In our multi-task framework, each question is paired with a corresponding key—forming a question-key pair—that encapsulates its unique characteristics and domain-specific attributes. This design enables a fine-grained classification of tasks, as each key serves as an identifier for the subdomain and the underlying problem-solving facet it represents. In our dataset, which is constructed from high-quality, authoritative UCLA prelim exams, each question is not only rigorously structured but also assigned a key that maps to distinct subdomains within Material Science. This approach allows for a combinatorial evaluation across various tasks and provides a nuanced understanding of the performance of large language models (LLMs). Moreover, the explicit association between questions and keys facilitates a higher resolution in assessment, offering insights into the model’s capability to address the subtle distinctions inherent in each category.
@@ D.3 AI Graders
Bias in grading arises when one arbitrator disproportionately favors certain ranks across repeated evaluations. For example, a grader may consistently assign lower (more critical) ranks, while another may tend to be more lenient. If such biases are not controlled, downstream evaluations may unfairly penalize or reward certain systems, obscuring the true effect of the Roundtable Policy. Consistency, on the other hand, reflects whether different graders produce comparable relative orderings. High consistency increases our confidence that the rankings reflect robust quality differences, while low consistency signals noise and arbitrariness. By quantifying both bias and consistency, we establish the validity of our evaluation methodology.
@@
We assess grader bias by visualizing the empirical rank distributions assigned by each arbitrator. Figure 3 presents kernel density estimates of the rank distributions for four graders across both the SingleTask and MultiTask evaluations. Systematic deviations in the distributions (e.g., skewed toward higher or lower ranks) indicate bias. To measure consistency, we compute pairwise Kendall’s Tau correlations between graders’ rankings. This non-parametric statistic captures the degree of ordinal agreement, with values close to indicating strong concordance, values near suggesting randomness, and negative values reflecting disagreement. The results are shown as inset heatmaps in Figure 3.
@@
In the SingleTask setting, we observe substantial variation in grader behavior. The rank distributions are uneven, with some graders systematically harsher or more generous than others. The Kendall’s Tau correlations confirm this observation, with moderate agreement values that indicate significant variability across graders. Such findings echo common phenomena in human peer review, where individual reviewers bring idiosyncratic standards and biases. By contrast, in the MultiTask setting, the distributions are more balanced, and the Kendall’s Tau values are consistently higher. This suggests that when grading across diverse tasks, individual biases are averaged out, leading to more stable and reliable inter-grader consensus.
@@
These analyses reinforce two important principles. First, reliance on a single grader is inherently risky due to potential bias; employing multiple graders and aggregating their judgments is essential for fairness. Second, broader task coverage increases reliability, as MultiTask grading mitigates the influence of any single arbitrator’s bias. Together, these results highlight the necessity of designing evaluation protocols that balance scalability with robustness, ensuring that AI grader-based assessments approximate the fairness and credibility traditionally associated with human expert panels.

### v2

@@ C.1 Dataset Annotation
We introduce Multi-Faceted Physical Reasoning, a heterogeneous benchmark designed to evaluate the scientific reasoning capabilities of LLMs in physical science and beyond. The benchmark spans six carefully curated sub-domains: (i) Macro Mechanics, (ii) Micro Mechanics, (iii) Optical & Microscopic, (iv) Semiconductor Physics, (v) Thermodynamical Systems, and (vi) Electrical & Electronic Properties. As visualized in Figure 10, these categories are balanced in both question count and token length, enabling systematic cross-domain evaluation of diverse reasoning skills.
@@
This hierarchical structure compels models to integrate reasoning across physical length scales—from atomic-level diffusion and microstructure evolution to bulk elasticity and device performance—while combining symbolic, conceptual, and numerical reasoning. Each sub-domain is paired with task-specific rubrics assessing both (a) factual accuracy and (b) explanatory depth. To further capture reasoning quality, all model outputs must include a chain-of-thought rationale, which is evaluated through two complementary methods: (i) the automatic Causal Reasoning Index (CRI), and (ii) a human-curated explanation rubric (details in Appendix C.4). These fine-grained metrics allow for robust measurement of model performance and principled ablation studies. Table 5 provides representative examples across four broader scientific areas—Biology, Geoscience, Mathematics, and Physical Science—highlighting domain-specific instruction design and diverse task objectives. Unlike prior benchmarks focused narrowly on short factual queries, our dataset emphasizes heterogeneous scientific tasks requiring numerical derivation, mechanistic explanation, and symbolic fluency.
Multi-Faceted Physical Reasoning.
@@
The physical-science questions are drawn from the 2006–2024 UCLA Ph.D. qualifying examinations. This provenance ensures originality, rigor, and relevance to authentic scientific reasoning. Each problem is annotated with: (i) a numeric ground-truth answer rounded to three significant digits, (ii) an exact LaTeX expression preserving symbolic fidelity, and (iii) a step-by-step expert-verified solution. All annotations are fully compliant with LaTeX and Markdown standards, ensuring reproducibility and integration with downstream evaluation pipelines.
@@
Subdomain Task Instruction More details Math Task 007: Math Janet’s ducks lay 16 eggs per day. She consumes 3 eggs for breakfast, uses 4 eggs for baking muffins, and sells the remaining eggs at $2 each. How much does she earn at the farmers’ market? Display strong arithmetic and logical sequencing skills to compute the final earnings. Geoscience Task 008: Geo. Archaeologists record the top layers first, followed by older, deeper layers to study ancient settlements according to the principle of superposition. Demonstrate the ability to interpret stratigraphic evidence and correctly apply the principle of superposition. Biology Task 009: Bio. Does tranexamic acid reduce desmopressin-induced hyperfibrinolysis? Link pharmacological interventions with clinical outcomes based on empirical data. Physical Science Macro Mechanics Task 001: Mech. 1 What is elasticity? What is strain energy? How do we make use of high elastic modulus materials? Explain fundamental mechanical properties and their practical applications in structural design. Micro Mechanics Task 002: Mech. 2 What are interstitial and substitutional diffusion? How do they differ from self-diffusion, and what drives these processes? Differentiate between diffusion mechanisms and explain the role of chemical potential gradients. Electrical & Electronic Task 003: EE What is the difference between homogeneous and heterogeneous nucleation? Demonstrate understanding of nucleation processes and how surfaces or impurities influence phase transitions. Optical & Microscopic Task 004: Opt. Using characterization techniques, how would you distinguish between a mixture of 90% NaCl + 10% LiCl and a mixture of 90% NaCl + 5% LiCl + 5% KCl? Interpret diffraction patterns or elemental analysis data to differentiate material compositions. Thermodynamical Task 005: Thermo. Explain the laws of thermodynamics both mathematically and conceptually. Articulate both the formal mathematical expressions and the underlying concepts of the thermodynamic laws. Semiconductor Task 006: Semi. What is Schrödinger’s equation? Use it to solve for a 1D wavefunction in a simple potential. Demonstrate a solid understanding of quantum mechanics by providing the correct form of Schrödinger’s equation and a step-by-step solution for a 1D potential problem.
@@ C.2 Dataset Comparison
| Geoscience(Deng et al., 2024) | 1395 | 109,159 |
| Biology(Jin et al., 2019) | 1000 | 92,866 |
| Math(Cobbe et al., 2021) | 1319 | 246,238 |
| Physical Science | 539 | 46,150 |
Table 6 benchmarks Multi-Faceted Physical Reasoning against a range of prominent scientific QA datasets. Whereas prior benchmarks often emphasize scale—providing thousands of short, loosely categorized questions—Multi-Faceted Physical Reasoning is designed for conceptual depth, compositional reasoning, and graduate-level rigor. For instance, the physical science subset averages tokens per question, more than double that of typical mathematics datasets, reflecting the structural and semantic complexity encountered in authentic scientific workflows.
@@ C.3 QK-Pair Overview
ScienceEval. Each round presents a batch of heterogeneous tasks spanning multiple scientific domains as shown in Figure 11. The AI Committee jointly evaluates player responses, testing the arbitration system’s ability to generalize across diverse input types, modalities, and difficulty levels.
ScienceNarrative. Each round focuses on a scientific proposal generation as shown in Figure 12, allowing the Committee to learn nuanced arbitration strategies within a fixed domain and structured output format.
Together, these protocols provide a comprehensive evaluation of arbitration dynamics under both breadth-oriented (ScienceEval) and depth-oriented (ScienceNarrative) conditions. Importantly, all evaluations are conducted without modifying the internal parameters of the player models, thereby preserving their pretrained competencies while enabling compositional reasoning through learned arbitration.
C.4 Multi-Faceted Physical Reasoning
Table 7 illustrates the mapping between several UCLA courses (from the prelim exams spanning 2006–2024) and our six defined subdomains in Multi-Faceted Physical Reasoning. The table indicates the number of questions per course and marks (✓) the corresponding subdomains covered by each course. And Table 8 Mapping from numeric course codes to official course names.
In ScienceEval, each question is paired with a corresponding key—forming a question-key pair—that encapsulates its unique characteristics and domain-specific attributes. This design enables a fine-grained classification of tasks, as each key serves as an identifier for the subdomain and the underlying problem-solving facet it represents. In our dataset, which is constructed from high-quality, authoritative UCLA prelim exams, each question is not only rigorously structured but also assigned a key that maps to distinct subdomains within Physical Science. This approach allows for a combinatorial evaluation across various tasks and provides a nuanced understanding of the performance of LLMs. Moreover, the explicit association between questions and keys facilitates a higher resolution in assessment, offering insights into the model’s capability to address the subtle distinctions inherent in each category.
@@ D.3 AI Graders
Bias in grading arises when one arbitrator disproportionately favors certain ranks across repeated evaluations. For example, a grader may consistently assign lower (more critical) ranks, while another may tend to be more lenient. If such biases are not controlled, downstream evaluations may unfairly penalize or reward certain systems, obscuring the true effect of Roundtable Policy. Consistency, on the other hand, reflects whether different graders produce comparable relative orderings. High consistency increases our confidence that the rankings reflect robust quality differences, while low consistency signals noise and arbitrariness. By quantifying both bias and consistency, we establish the validity of our evaluation methodology.
@@
We assess grader bias by visualizing the empirical rank distributions assigned by each arbitrator. Figure 7 presents kernel density estimates of the rank distributions for four graders across both the ScienceEval and ScienceNarrative evaluations. Systematic deviations in the distributions (e.g., skewed toward higher or lower ranks) indicate bias. To measure consistency, we compute pairwise Kendall’s Tau correlations between graders’ rankings. This non-parametric statistic captures the degree of ordinal agreement, with values close to indicating strong concordance, values near suggesting randomness, and negative values reflecting disagreement. The results are shown as inset heatmaps in Figure 7.
@@
In the ScienceNarrative, we observe substantial variation in grader behavior. The rank distributions are uneven, with some graders systematically harsher or more generous than others. The Kendall’s Tau correlations confirm this observation, with moderate agreement values that indicate significant variability across graders. Such findings echo common phenomena in human peer review, where individual reviewers bring idiosyncratic standards and biases. By contrast, in the ScienceEval, the distributions are more balanced, and the Kendall’s Tau values are consistently higher. This suggests that when grading across diverse tasks, individual biases are averaged out, leading to more stable and reliable inter-grader consensus.
@@
These analyses reinforce two important principles. First, reliance on a single grader is inherently risky due to potential bias; employing multiple graders and aggregating their judgments is essential for fairness. Second, broader task coverage increases reliability, as ScienceEval grading mitigates the influence of any single arbitrator’s bias. Together, these results highlight the necessity of designing evaluation protocols that balance scalability with robustness, ensuring that AI grader-based assessments approximate the fairness and credibility traditionally associated with human expert panels.
Appendix E Multi-Agent Baselines
Majority Vote (Wang et al., 2023). In the majority vote baseline, multiple agents independently generate candidate responses to the same query. The final output is selected by simple majority voting over the agents’ responses. No historical information or agent-specific weighting is used; all agents contribute equally at inference time.
Weighted Vote (Taubenfeld et al., 2025). Weighted vote extends majority voting by assigning different weights during aggregation. In our implementation, weights are determined by self-reported confidence scores provided by the agents. The final prediction is obtained through confidence-weighted aggregation rather than uniform voting.
Debate (Du et al., 2024). Debate is an interaction-based multi-agent framework in which agents iteratively exchange arguments to refine their responses. Agents begin by independently generating initial answers, followed by multiple rounds of critique and revision. After the final debate round (three rounds in our implementation), the system outputs one agent’s response, corresponding to the final updated solution.
Debate w Judge (Liang et al., 2024). Debate w Judge augments the standard debate framework by introducing an explicit adjudicator agent. Following several rounds of debate (three rounds in our implementation), the adjudicator reviews all final responses and selects the most convincing answer. To ensure architectural consistency across methods, we employ the same aggregator model used in Roundtable Policy as the judge. The adjudicator is applied only at the final decision stage and does not intervene during the debate process itself.
Debate w Confidence (Chen et al., 2024). Debate w Confidence further extends debate by incorporating agent confidence signals into the aggregation process. The framework proceeds in three phases. First, each agent independently generates an initial response and accompanying explanation. Second, agents engage in multiple rounds of discussion (three rounds in our implementation), updating their responses based on information exchanged with peers. As both the ScienceEval and ScienceNarrative settings involve open-ended reasoning rather than discrete choices, the grouping mechanism is instantiated as concatenation of peer responses. Finally, after each discussion round, the system produces an output via confidence-weighted voting over agent responses.
Mixture-of-Agents (Wang et al., 2025). Mixture-of-Agents is a layered multi-agent aggregation framework that exploits the observed collaborativeness of large language models, whereby response quality can improve when conditioned on outputs from other agents. In our implementation, we adopt a three-layer architecture with nine agents per layer. The same aggregator model as in Roundtable Policy is used for the final aggregation stage to maintain consistency across methods.